OpenStudy (mellamokatie):

When solving a system of equations, how do you determine which method to use? As in substitution and elimination and graphing. I need help with examples and definitions. I don't understand what it is and I have a written response question. Thanks!

OpenStudy (photon336):

well, they both lead you to the correct answer so I think substitution would be a good method to use followed by maybe graphing to check the answer you get.

OpenStudy (mellamokatie):

No, you don't understand I have to explain them all with examples....

OpenStudy (mellamokatie):

@sweetburger

OpenStudy (mellamokatie):

@sammixboo |

Vocaloid (vocaloid):

1. graphing graphing just means you graph all the equations in the system, and identify where the graphs intersect. the intersection points are your solutions

OpenStudy (mellamokatie):

Thank you!!

Vocaloid (vocaloid):

2. substitution substitution means solving for one variable in one of the equations, then substituting it into another equation ex. say I have x + y = 2 and 2y = 1 I can solve the second equation for y (y = 1/2) and substitute this y value into the first equation x + y = 2 x + (1/2) = 2 x = 3/2 then I can solve for y by using either of the original equations

Vocaloid (vocaloid):

3. elimination rewrite both equations in such a way that one or more variables can be eliminated through addition or subtraction of the equations

Vocaloid (vocaloid):

ex: x + 2 = y -2x + y = 3 we can rewrite both of them in standard form to get x + y = -2 -2x + y = 3 then we can multiply the first equation by 2 to get 2x + 2y = -4 -2x + y = 3 and add them together to eliminate x 2x - 2x = 0 2y + y = 3y -4 + 3 = -1 therefore 3y = -1 y = -1/3 then to find x we just substitute -1/3 for y in one of the original equations and solve for x

OpenStudy (mellamokatie):

Wow thanks for all this work!! Thank you!! Wish I could give you more than just a medal!

Vocaloid (vocaloid):

oh i almost forgot

Vocaloid (vocaloid):

it asks how you decide which method to use

OpenStudy (nincompoop):

all this work

OpenStudy (nincompoop):

she will get an A and thanks to you

OpenStudy (mellamokatie):

Yep agreed !!!

Vocaloid (vocaloid):

i think i gave out more help than I was supposed to

Vocaloid (vocaloid):

but I hope you get the idea, just don't copy it word for word

OpenStudy (mellamokatie):

Yeah I changed the wording a bit I will show you when I am done.